OpenStudy (anonymous):

What is wrong with the solution? x(x+3)=9 x=9 or x+3=9 x=6 So, the roots are 9 and 6.

OpenStudy (candyme):

Ok, to solve by that rule the right side of the equation has to be zero. You would have to multiply it out, combine like terms, and then solve.

OpenStudy (anonymous):

multiply 9 to ?

OpenStudy (candyme):

Multiply out the x(x+3). x^2+3x=9 Subtract 9 from both sides x^2+3x-9=0

OpenStudy (anonymous):

ohh i get it now thank you :)

OpenStudy (candyme):

Since it doesn't factor you have to use the quadratic formula

OpenStudy (anonymous):

so im going to factor x^2 + 3x - 9 = 0 ?

OpenStudy (anonymous):

ohh idk how to use the quadratic formula :((((

OpenStudy (candyme):

Ok, let me switch to my computer so I can draw this better

OpenStudy (candyme):

So the quadratic formula is\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a}\] and in the equation the coefficients of your problem are the a, b, and c. \[1x^2+3x-9=0\]1 is your a 3 is your b -9 is your c

OpenStudy (candyme):

Now, since it's a plus or minus sign (so you get both roots) you will have 2 equations The first one\[\frac{ -1+\sqrt{(3^2)-(4)(1)(-9)} }{ (2)(1) }\] And the second one\[\frac{ -1-\sqrt{(3^)-(4)(1)(-9)} }{(2)(1) }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

whats next

OpenStudy (candyme):

If you solve those, you get your roots.

OpenStudy (candyme):

\[\frac{ -1+\sqrt{45} }{2 }=2.854101966\] is the first equation. I'm not sure whether you need it completely simplified or in radical form. The second equation is\[\frac{ -1-\sqrt{45} }{ 2 }=-3.854101966\]

OpenStudy (anonymous):

thank you so much! i already fan(ned) :)

OpenStudy (candyme):

Aww Thanks! Glad I could help:)

OpenStudy (ashleyisakitty):

Great work @candyme!

OpenStudy (anonymous):

uhhh is the second equation 3squared too?

OpenStudy (candyme):

Yes, I didn't even notice the typo